# SPECIAL RELATIVITY

### v9

@@ 8.6 Absent of the Aberration of Light from the Earth-based Laser Source
Let us analyze the aberration of light transmitted from the laser resonator which is at rest in the accelerated frame. When the measuring data is analyzed, the accelerated observer finds that the deviation of the energy transport direction is absent. It could be said that the crossed term in the Langevin metric, which generates aberration, cancels during the (round-trip-to-round-trip) evolution of the radiation in the optical resonator.
We would like to emphasize a very important difference between laser and incoherent source. The radiation of incoherent source clearly depends on an initial conditions and there is influence of the mixture of positions and time which is a result of pseudo-gravity experienced by the accelerated observer. In contrast, when we are dealing with optical resonator the (3-space) boundary conditions define the energy propagation direction.
Now let us consider another problem in optical resonator operation that we should briefly discuss. This has to do with the optical experiments of second order. It is important to remark that a close examination of all experiments inside the uniformly moving accelerated $S_{n}$ frame, however, shows that all optical phenomena in the optical resonator appeared to be independent of the uniform motion relative to the fixed stars.
Imaging that there are two identical laser sources. We set up the lasers such that light is travels between mirrors parallel to the $z$-axis. Let us consider the case when the first source is at rest in an inertial frame and the second source is accelerated from rest up to velocity $v$ along the $x$-axis (i.e. transverse to the direction of optical beam). Suppose that an observer in the inertial frame performs a measurement of the frequency $\omega_{m}=m\pi c/l$ of the $TEM_{00}$ mode. In order to measure the frequency, the observer has to specify frequency standard. Suppose he took an ordinary atom, which had natural frequency $\omega_{o}$ at rest. Let us consider the opposite situation. Suppose that an observer in the accelerated frame performs a measurement of the $TEM_{00}$ mode. It is also assumed that the same atom at rest in the accelerated frame is used as frequency standard. In the case of Langevin metric Eq.(15), the speed of light propagated between the mirrors in the transverse $z_{n}$-axis direction cannot be equal $c$ anymore. Using Eq.(15) we obtain $(dz_{n}/dt_{n})^{2}=c^{2}(1-v^{2}/c^{2})$. Therefore, the frequency of the $TEM_{00}$ mode is $\omega_{m}=m\pi c\sqrt{1-v^{2}/c^{2}}/l$. Because $t_{n}=t$ (note that coordinate clock in the accelerated frame , although fixed in this frame , read the same time as the clocks in the inertial frame) and $\omega_{m}=m\pi c/l$ has the physical meaning of the frequency of the dominate mode at rest in the inertial frame, this implies that the frequency of the $TEM_{00}$ mode at rest in the accelerated frame is red shifted compared to $\omega_{m}$ in the original inertial frame. Taking into account the Langevin metric we obtain the time dilation of a physical clock at rest in the accelerated frame as compared to physical clock in the inertial frame. In the next chapter we shall discuss the asymmetry in the experiments involved the measurement of quantities of the second order in $v/c$.
An observer on the accelerated frame cannot observe the time dilation effect using optical resonator. In order to detect the red shift of the $TEM_{00}$ mode at rest in the accelerated frame, it is obvious that some reference frequency is needed. The point is that the natural frequency of the atom observed in the accelerated frame is $\omega_{0}\sqrt{1-v^{2}/c^{2}}$. We have found that the accelerated observer will measure the same mode frequency. It is not hard to understand this result. It is clear that an observer performs the measurement of standing wave mode of the optical resonator. If we analyze the geometry of the situation, we find that from the standing wave measurements in the optical resonator we can only extract information about two-way speed of light. In other words, the measurement of the two-way speed of light is universal. Although the optical cavity we have been describing seems to be quite different from a simple electromagnetic model for the light-clock of special relativity, the two systems are, of course closely related. The discussion of the behavior of an accelerated light clock is presented in the next chapter.
8.7 Bibliography and Notes
1. The dedicated STEM with high stability was made [39]. It was developed based on a 200-kV commercial instrument. The electron source was installed in the anti-seismic room. Temperature fluctuations in the room were within 0.2 ${}^{o}C$. The magnetic field in the room was less than $0.2mG$ near the source column. To prevent thermal drift, a shield box was placed over the high-voltage generator. The column was covered with rubber sheets to reduce the temperature fluctuations. The image drift measured during a period of 61 minutes was $\simeq 6~{}\mathrm{nm}$. The drift speed and direction during this period were approximately constant. The average drift speed was 0.12 nm/min. It is substantially smaller than that conventional commercial microscope, in which the drift speed is about 2 nm/min.
2. The rotation of the earth on its axis produces aberration in an amount large enough to be taken into account in precise observation work using electron microscope as a particle source. The Wigner rotation angle parameter (constant of diurnal aberration) defines the scale of the aberration increment, and is equal to $\theta_{w}=(1-1/1/\gamma_{p})v_{e}/V\simeq 0.6\mu\mathrm{rad}$. Here $\gamma_{p}\simeq 1.39$, $V/c=0.7$, $v_{e}/c\simeq 1,55\mu\mathrm{rad}$ where $v_{e}=0.45$km/s is the linear velocity of the earth rotation at equator.
3. Electron-optical system of the STEM contains an electron gun and several magnetic lenses, stacked vertically to form a lens column. The illumination system of the instrument comprises the electron gun, together with two condenser lenses that focus the electrons onto the spacemen. Electrons entering the lens column appear to come from a virtual source. The first condenser lens (C1) is a strong magnetic lens, with a focal length $f_{1}$ of typically 2 mm. Using a virtual electron source as its object, C1 produces a real image. Because the lens is located $D_{1}=15$ cm bellow the object (virtual source), its object distance is 15 cm $\gg f_{1}$ and the image distance $\simeq f_{1}$. The second condenser lens (C2) is a weak magnetic lens ($f_{2}\simeq 10$ cm) that provides little magnification but allows the diameter of illumination at the specimen to varied continuously over a wide range. Distance between lens centers typically 10 cm. Spacemen is located $D_{3}=25$ cm bellow the C2.
4. Experimental results (see Fig. 44 p.167 in [39]) shows that the image shift is quite close to the theoretical prediction $\simeq 10~{}\mathrm{nm}$. It is important to note that we used the typical focal length (2 mm) in our estimations. It is also assumed that the second condenser lens provides no magnification ($M\simeq 1$). The details of the optical setup may influence the image shift.
5. The original optical source (such as a star or an incandescent lamp) consists of an extended collection of independent radiators. Accordingly, it is of some special interest to know precisely how mutual intensity propagates away from an incoherent source. The character of the mutual intensity function produced by an incoherent source is fully described by the Van Cittert-Zernike theorem [34].
6. In all cases of practical interest, telescopes are situated in the far zone of the distant stars. If the far zone approximation $z>2d^{2}/\lambda$ is satisfied, then the Van Cittert-Zernike theorem takes its simplest form. For example, the coherence area of the light emitted by the circular incoherent source of diameter $d$ at distance $z$ is $A_{c}=4\lambda^{2}z^{2}/(\pi d^{2})$. The maximal coherent area of light observed on the earth’s surface from a nearest star has diameter of about 10 m. In other words, any star can be considered as a point source and a star image is actually a point spread function in the image plane of a telescope. The star consists of an extended collection of independent ”point” sources. The order of magnitude of the dimensions of the elementary statistically independent source is about $\lambda$. It is obvious that the star diameter $d$ is much larger than a telescope aperture $D$ and that far zone condition $z>D^{2}/\lambda$ for elementary ”point” source is automatically satisfied. Our scaling model is based on the far zone approximation. We have only to change the distance and the source diameter. With the limitation of the small far zone parameters ($d^{2}/(z\lambda)<1$, $D^{2}/(z\lambda)<1$) assumption, the proposed scaling method gives $d\ll D$.
7. It may be possible to use the masked spatially incoherent UV light source. One could use the light source onto (micron diameter) hole of specially designed mask. The mask should be placed at the distance $l<<d$, where $d$ is the transverse size of the spatially incoherent UV light source.
8. The important question is that of light detection. The small image size form rather sever restriction on the use of the CCD as a detector. In fact, for the case considered numerically image size is near 2 $\mu$m. One possibility to overcome this practical difficulty is the photon detection with an MCP (micro channel plate photo multiplier) detector. It may be possible to use a (commercially available) precision 2D translation stage with a resolution of 0.5 $\mu$m to controlling the motion of a hole in the opaque screen. The aperture diameter in this example may be set to 1 micron or smaller. The screen is fixed horizontally on the translation stage which is placed in the image plane. The MCP-PMT may realize the measurement of (transmitted through a hole) photons. The proposed design has the advantage of device compactness. The height of our proposed optical configuration is about one meter.
9. We have presented an example of introducing an approximate boundary conditions. This approximate boundary conditions were derived by Leontovich. They are valid on the surface of a material having a large value of the modulus of the refractive index. An extended theoretical treatment of Leontovich’s boundary conditions appears in the monograph [40].
10. Analytical method is based on the introduction of particular, complex boundary condition, which are called impedance boundary conditions [41]. Once these boundary conditions for the field are formulated on the virtual side surface of the open plane resonator, the original problem, which is an open resonator excitation, is simplified to the that of a closed resonator.
11. $\beta_{0}$ turns out to be related with one of the most famous mathematical functions, the Riemann zeta function $\zeta$. In fact, it is given by $-\zeta(1/2)/\sqrt{\pi}$, see [41].
@@ 9 Special Relativity and the Reciprocal Symmetry
In this chapter we will collect a number of useful facts concerning the space-time measurements made by different observers moving with respect to each other. The situation with measurements in special relativity is not symmetrical with respect to the change of the reference frames. The fact that in the real process of transmission to an accelerated reference frame (i.e in the process of an observer accelerating with respect to the fixed stars) the observer will experience the pseudo-gravitational force, is not considered in textbooks.
@@ 9.1 Pseudo-gravitation Field and the Langevin Metric
This section concentrates on the nature of special relativistic effects. Dynamics, based on the field equations, is usually hidden in the language of the relativistic kinematics. The metric structure of space-time is only an interpretation of the behavior of the dynamical matter fields in the view of different observers. It is worth remarking that the absent of a dynamical explanation for the Langevin metric in special relativity has disturbed some physicists. A good way to think of the asymmetry between the inertial and accelerated frames is to regard it as a result of pseudo-gravity experienced by the accelerated observer.
Langevin metric associated with the transformation from the inertial coordinate system to the accelerated (with respect to the fixed stars) coordinate system may be regarded as a result of the action of certain force. This problem formulated within the framework of general relativity. Note that it is not correct to say that this is explanation of the effects discussed in the last four chapters. One of the aims in this section is to show that one can describe the non-inertial frames in the pseudo-Euclidean space-time using the language of general relativity.
@@
When the system $S_{n}$ starts moving with constant velocity it will be inertial frame of reference. At first glance, we have the symmetry (reciprocity) between the inertial frame $S$ and inertial frame $S_{n}$. According to principle of covariance (i.e. principle of reciprocity), accelerated (with respect to the fixed stars) observer should obtain Minkowski metric $ds^{2}=c^{2}dt_{n}^{2}-dx_{n}^{2}$. This metric is reciprocal to the metric of the inertial observer $ds^{2}=c^{2}dt^{2}-dx^{2}$. Where does the asymmetry comes from? Equivalence principle analysis does not need any tricks to account for the asymmetry between an inertial frame and an accelerated (with respect to the fixed stars) frame. A resolution of the asymmetry paradox identifies a pseudo-gravitational potential within the system $S_{n}$ as the agency of asymmetry.
@@
We can treat a uniformly accelerating frame as if it were an inertial frame with the addition of a uniform pseudo-gravitational field. By a ”pseudo-gravitational field”, we mean an apparent field (not a real gravitational field) that acts on all objects proportionally to their mass; by ”uniform” we mean that the force felt by each object is independent of its position. This is basic content of equivalence principle. The principle of equivalence can be applied to solve non-inertial kinematic problems with dynamics methods.
@@
In order to keep the mathematical complexity of the discussion to a minimum, we will describe the effects by working only up to the second order in $v/c$. Suppose that each reference point in the system $S$ has a constant acceleration $g=-|g|$ in the negative direction along the $x_{n}$- axis during the time $T=v/|g|$. In other words, the system $S$ is accelerating (freely falling) in a pseudo-gravitational field. It is now immediately clear that the pseudo-gravitational acceleration is simply equal to the gradient of the scalar potential: $g=-\partial\phi/\partial x_{n}$. A simple calculation gives $x_{n}=gT^{2}/2+x$ at $t=T$. When the system $S$ starts moving with constant velocity the pseudo-gravitational field is zero. For the pseudo-gravitational potential difference between the two frames we get after some calculation $\phi=(\partial\phi/\partial x_{n})gT^{2}/2=-v^{2}/2$. It is assumed that $\phi=0$ at $x_{n}=x$.
@@
The equivalence principle analysis of the accelerated frame does not use any real gravitational field, and so does not use the general relativity. Nevertheless, what general relativity does say about real gravitational fields does hold in the restrictive sense for pseudo-gravitational fields. One thing we need here is that time runs slower as we descend into the potential well of a pseudo force field. We can use that fact to our advantage when analyzing the metric in the non-inertial frame.
@@
The equivalence principle implies that gravity can shift the frequency of an electromagnetic wave, and cause clocks to run slow. The light frequency increases with increase absolute value of the potential of the gravitational field. If a pulse of light emitted at a point where the gravitational potential is $\phi_{1}$, has at that point the frequency $\omega_{1}$, then upon arriving at a point where the potential is $\phi_{2}$, it will have a frequency measured in units of the proper time at that point equal to $\omega_{2}=\omega_{1}[1+(\phi_{1}-\phi_{2})/c^{2}]$.
@@
Frequency is proportional to the inverse of the local proper time rate; the gravitational frequency shift formula can be converted to a time dilation formula. The rate of a coordinate clock thus depends on the gravitational potential at the place where the clock is situated. The relation between coordinate time and physical (proper) time can be written in the form $dt^{(p)}=dt_{n}(1+\phi/c^{2})$. Thus physical time elapses the more slowly the smaller the gravitation potential at a given point in a space, i.e. the lager the absolute value. If of two identical clocks is placed in a gravitation field for some time, the clock which has been in the field will thereafter appear to be slow. For clock at rest in the accelerated frame we have $dt^{(p)}=dt_{n}(1+\phi/c^{2})=dt_{n}[1-v^{2}/(2c^{2})]$. The fact that the inertial observer always turns out to have aged more than the non-inertial one, in the equivalence principle analysis is elementary consequence of the slowly down of the rate of the processes in the pseudo-gravitational potential, which can be shown by means of simple algebraic calculations.
@@
Furthermore it is important to point out that the phenomenon of the relativity o simultaneity can be understood in terms of dynamical consideration. The time-offset relation can be interpreted as the accumulated time difference between two spatially separated clocks because of pseudo-gravity experienced by the accelerated observer. This suggests that the clock at higher gravitational potential (placed along the direction of acceleration) runs faster. For the pseudo-gravitational potential difference between two points along the $x_{n}$-axis we get $\Delta\phi=\phi_{1}-\phi_{2}=(\partial\phi/\partial x_{n})[x_{n}(1)-x_{n}(2)]$. When the system $S$ starts moving with constant velocity the gradient of potential in the system $S_{n}$ is zero. For accumulated time difference between two spatially separated clocks we have $t^{(p)}(1)-t^{(p)}(2)=-g[x_{n}(1)-x_{n}(2)]v/(c^{2}|g|)=v[x_{n}(1)-x_{n}(2)]/c^{2}$.
@@
The next thing we must investigate is the origin of the Lorentz deformation. Here we may point out that the Lorentz deformation (changes of the relative position in the $x_{n}$ direction) is closely related to the shift in the moments of the acceleration. With the help of time-offset relation, integrating with respect to the infinitesimal value $dv$ we find $dl_{n}=dx_{n}[1+v^{2}/(2c^{2})]$, which determines the spatial geometry in the accelerated frame. This equation relates the length $dl=dx_{n}$ of the rod in the inertial frame $S$ to the length $dl_{n}$ in the accelerated frame $S_{n}$. Our analysis based on using standard measuring rods in an accelerated frame to measure it geometrical properties shows that coordinate distance $dx_{n}$ has length $dx_{n}[1+v^{2}/(2c^{2})]$. This is a purely kinematic effect, involving no forces. We come to the conclusion that the various relativistic effects are not independent of each other, but that these effects can all be understood through a general equivalence principle.
@@
We have now all quantities we wanted. Let us put them all together into the expression for the new space-time variables. The new independent variables $(x_{n}^{\mathrm{(p)}},t_{n}^{\mathrm{(p)}})$ can be expressed in terms of the old independent variables $(x_{n},t_{n})$: $ct_{n}^{\mathrm{(p)}}=t_{n}[1-v^{2}/(2c^{2})]-(vx_{n}/c^{2})[1+v^{2}/(2c^{2})]$, $x_{n}^{\mathrm{(p)}}=x_{n}[1+v^{2}/(2c^{2})]$. Above we made a simplification in calculating time and space transformations by considering only low velocities. The physical principles that produced the space-time transformation were, however, made clear. The next question is: What is the transformation in the case of an arbitrary velocity? That is easy: the answer is $t_{n}^{\mathrm{(p)}}=t_{n}\sqrt{1-v^{2}/c^{2}}-(vx_{n}/c^{2})/\sqrt{1-v^{2}/c^{2}}$, $x_{n}^{\mathrm{(p)}}=x_{n}/\sqrt{1-v^{2}/c^{2}}$. Quantity $dt_{n}^{\mathrm{(p)}}$ characterizes physical time, which is independent on the choice of coordinate time. Let us note that $t_{n}$ in the expression for $t_{n}^{\mathrm{(p)}}$ is the coordinate time for the accelerated frame. Physical time determines the flow of time in a physical process. The $|dx_{n}^{\mathrm{(p)}}|$ is nothing, but the physical distance between two points of 3-space. In new physical coordinates we obtain interval in the following (Langevin) form $ds^{2}=(cdt_{n}^{\mathrm{(p)}})^{2}$ $(dx_{n}^{\mathrm{(p)}})^{2}$ $=(1-v^{2}/c^{2})c^{2}dt^{2}_{n}-2vdx_{n}dt_{n}-dx_{n}^{2}$.
@@
The previous derivation of the Langevin metric in the accelerated frame from pseudo-gravitation field includes one delicate point. Metric applies to physical laws, not to physical facts. We interpret the Langevin metric to mean that the law of electrodynamics is expressed by equation that have the form Eq.(16) in the accelerated frame. By a physical facts in this context we mean, for example, the aberration of light radiated by a single ”plane wave” emitter in the accelerated frame. The electrodynamics equation needs to be integrated with initial condition for the radiation wavefront. After the inverse Galilean boost, the coordinates transform as $x_{n}=x-vt,t_{n}=t$, so that the wavefront of the emitted light beam is perpendicular to the vertical direction $z_{n}$ after the acceleration, Fig. 15. The Langevin metric, together with this initial condition, describes the aberration of light effect in the accelerated frame.
Although both observers consider themselves as at rest, they both predict that the Maxwell’s equations continue to hold in the inertial frame and the Minkowski metric is not applicable in the accelerated frame. But observers have totally different explanation for this difference of metrics. Inertial observer says that physical processes in the accelerated frame describes by the Langevin metric due to the active Galilean boost. In other words, the analysis of physical phenomena in non-inertial frames of reference can be described in an inertial frame within standard pseudo-Euclidean geometry of space-time. Accelerated observer, on the other hand, says that Langevin metric in his frame can be understood in terms of dynamical consideration.

### v10

@@ 11.2 Absent of the Aberration of Light from the Earth-Based Laser Source
Let us now analyze the aberration of light transmitted from the laser resonator which is at rest in the accelerated frame. When the measuring data is analyzed, the accelerated observer finds that the deviation of the energy transport direction is absent, Fig. 43. It could be said that the crossed term in the Langevin metric, which generates aberration, cancels during the (round-trip-to-round-trip) evolution of the radiation in the optical resonator.
Figure 43: Aberration of light in an accelerated frame of reference. It is assumed that the accelerated laser operates in the steady-state regime.
Figure 44: Aberration of light from a ”plane wave” emitter in an accelerated frame of reference.
According to the difference between the ”plane wave” emitter and the laser source, there is a remarkable prediction in the theory of the aberration of light. Namely, if the ”plane wave” emitter was at rest relative to the fixed stars and this source started from rest to motion relative to the fixed stars, then the apparent angular position of the ”plane wave” emitter seen in the accelerated frame would jump by angle $-v/c$, Fig. 44. To understand the aberration of light, we will have to use the special theory of relativity and not be satisfied with classical theory.
How can such a difference between a ”plane wave” emitter and a laser source come about? Let us first ask: ”What happens to accelerated optical resonator?” When we were dealing with optical resonators we used the electrodynamics boundary conditions in 3-space, and they defined the energy propagation direction along the optical axis (i.e. perpendicular to the mirror surface). According to special relativity, an acceleration does not change the direction of the optical axis (i.e. mirrors orientation) with respect to the accelerated frame (3-space) axes. An observer in the rest frame of a laser source finds that the direction of the energy transport is independent of the motion with respect to the fixed stars, Fig. 43.
According to special relativity, accelerated and inertial observers have common absolute space-time but have different metrics and, consequently, different 3-spaces. When we are dealing with optical resonator the (3-space) boundary conditions define the energy propagation direction. In contrast, the radiation of an incoherent source depends on initial conditions and there is an influence of the mixture of positions and time which is a result of pseudo-gravity experienced by the accelerated observer. The solution is presented in terms of the absolute time coordinate, hence the wavefront of the emitted light beam is perpendicular to the direction $z_{n}$ after the acceleration. Using this initial condition and the Langevin metric in a frame of reference attending acceleration enables us to explain the aberration of light from a ”plane wave” emitter, Fig. 44.
11.3 A Moving Laser Source in an Inertial Frame of Reference
Let us consider the radiation emitted by the accelerated laser source in the inertial frame. The aberration of light effect can be described within the standard special relativity taking advantage of the relativity of simultaneity. When one has a transversely (i.e. parallel to the mirrors) moving laser there is a deviation of the energy transport for light transmitted from the optical resonator. In the first order approximation in $v/c$, the inertial observer would find that angular displacement is $\theta_{a}=2v/c$, Fig. 45.
In regard to light aberration one should differentiate between that from the ”plane wave” emitter and that from the laser source. Suppose we have a ”plane wave” emitter. We consider the case when the ”plane wave” emitter is accelerated from rest up to velocity $v$ in the direction perpendicular to its optical axis. Suppose that an observer, which is at rest with respect to the inertial frame of reference performs the direction of the light transport measurement. Then how does the light beam from the moving ”plane wave” emitter look? The inertial observer would find that angular displacement is equal to $\theta_{a}=v/c$, Fig. 46.
Figure 45: Aberration of light from a moving laser source in the initial inertial frame of reference. The motion of the laser resonator is parallel to the mirrors. It is assumed that the accelerated laser operates in the steady-state regime.
Figure 46: Aberration of light from a moving ”plane wave” emitter in the initial inertial frame of reference.
The first thing we would say about inertial frame measurements is that there is an intuitively plausible way to understand the aberration of light from a moving ”plane wave” emitter. An elementary explanation of this effect is well-known. This phenomenon is fully understandable in terms of the transformation of velocities between different reference frames. The aberration of light in the initial inertial frame can also be easily explained on the basis of the corpuscular theory of light. This is plausible if one keeps in mind that a light signal represents a certain amount of electromagnetic energy. Energy, like mass, is a quantity that is conserved, so that a light signal resembles, in many aspects, material particles. Therefore, we should expect that group velocities of light signals obey the same addition theorem for particle velocities. A closer treatment based on the wave theory of light confirms this expectation.
In the case of the moving laser source, intuition would seem to tell us that the aberration increment would be equal to $\theta_{a}=v/c$. But the special relativity says that there is an extra factor 2. The commonly accepted derivation of the aberration of light effect does not account for the mixture of positions and time in the accelerated frame. It is incorrectly assumed that accelerated and inertial observers have a common 3-space. In contrast to this, as viewed from the initial inertial frame, the coordinate axes of the accelerated frame are rotated through the angle $v/c$ with respect to the coordinate axes of the inertial frame.
@@ 12.1 Pseudo-gravitation Field and the Langevin Metric
This chapter concentrates on the nature of special relativistic effects. Dynamics, based on the field equations, is usually hidden in the language of relativistic kinematics. The metric structure of space-time is only an interpretation of the behavior of the dynamical matter fields in the view of different observers. It is worth remarking that the absence of a dynamical explanation for the Langevin metric in special relativity has disturbed some physicists. A good way to think of the asymmetry between the inertial and accelerated frames is to regard it as a result of pseudo-gravity experienced by the accelerated observer.
Langevin metric associated with the transformation from the inertial coordinate system to the accelerated (with respect to the fixed stars) coordinate system may be regarded as a result of the action of a certain force. This problem is formulated within the framework of general relativity. Note that it is not correct to say that this is an explanation of the effects discussed in the last chapters. One of the aims of this chapter is to show that one can describe the non-inertial frames in the pseudo-Euclidean space-time using the language of general relativity.
@@
When the system $S_{n}$ starts moving with constant velocity it will be the inertial frame of reference. At first glance, we have the symmetry (reciprocity) between the inertial frame $S$ and the inertial frame $S_{n}$. According to the principle of covariance (i.e. principle of reciprocity), accelerated (with respect to the fixed stars) observer should obtain Minkowski metric $ds^{2}=c^{2}dt_{n}^{2}-dx_{n}^{2}-dy_{n}^{2}-dz_{n}^{2}$. This metric is reciprocal to the metric of the inertial observer $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$. Where does the asymmetry come from? Equivalence principle analysis does not need any tricks to account for the asymmetry between an inertial frame and an accelerated (with respect to the fixed stars) frame. A resolution of the asymmetry paradox identifies a pseudo-gravitational potential within the system $S_{n}$ as the agency of asymmetry.
@@
We can treat a uniformly accelerating frame as if it were an inertial frame with the addition of a uniform pseudo-gravitational field. By a ”pseudo-gravitational field”, we mean an apparent field (not a real gravitational field) that acts on all objects proportionally to their mass; by ”uniform” we mean that the force felt by each object is independent of its position. This is the basic content of the equivalence principle. The principle of equivalence can be applied to solve non-inertial kinematic problems with dynamics methods.
@@
In order to keep the mathematical complexity of the discussion to a minimum, we will describe the effects by working only up to the second order in $v/c$. Suppose that each reference point in the system $S$ has a constant acceleration $g=-|g|$ in the negative direction along the $x_{n}$- axis during the time $T=v/|g|$. In other words, the system $S$ is accelerating (freely falling) in a pseudo-gravitational field. It is now immediately clear that the pseudo-gravitational acceleration is simply equal to the gradient of the scalar potential: $g=-\partial\phi/\partial x_{n}$. A simple calculation gives $x_{n}=gT^{2}/2+x$ at $t=T$. When the system $S$ starts moving with constant velocity the pseudo-gravitational field is zero. For the pseudo-gravitational potential difference between the two frames, we get after some calculation $\phi=(\partial\phi/\partial x_{n})gT^{2}/2=-v^{2}/2$. It is assumed that $\phi=0$ at $x_{n}=x$.
@@
The equivalence principle analysis of the accelerated frame does not use any real gravitational field, and so does not use the general relativity. Nevertheless, what general relativity does say about real gravitational fields does hold in the restrictive sense for pseudo-gravitational fields. One thing we need here is that time runs slower as we descend into the potential well of a pseudo-force field. We can use that fact to our advantage when analyzing the metric in the non-inertial frame.
@@
The equivalence principle implies that gravity can shift the frequency of an electromagnetic wave, and cause clocks to run slow. The light frequency increases with the increase absolute value of the potential of the gravitational field. If a pulse of light emitted at a point where the gravitational potential is $\phi_{1}$, has at that point the frequency $\omega_{1}$, then upon arriving at a point where the potential is $\phi_{2}$, it will have a frequency measured in units of the proper time at that point equal to $\omega_{2}=\omega_{1}[1+(\phi_{1}-\phi_{2})/c^{2}]$.
@@
Frequency is proportional to the inverse of the local proper time rate; the gravitational frequency shift formula can be converted to a time dilation formula. The rate of a coordinate clock thus depends on the gravitational potential at the place where the clock is situated. The relation between coordinate time and physical (proper) time can be written in the form $dt^{(p)}=dt_{n}(1+\phi/c^{2})$. Thus physical time elapses the more slowly the smaller the gravitation potential at a given point in a space, i.e. the larger the absolute value. If two identical clocks are placed in a gravitation field for some time, the clock that has been in the field will thereafter appear to be slow. For clock at rest in the accelerated frame we have $dt^{(p)}=dt_{n}(1+\phi/c^{2})=dt_{n}[1-v^{2}/(2c^{2})]$. The fact that the inertial observer always turns out to have aged more than the non-inertial one, in the equivalence principle analysis is an elementary consequence of the slowdown of the rate of the processes in the pseudo-gravitational potential, which can be shown using simple algebraic calculations.
@@
Furthermore, it is important to point out that the phenomenon of the relativity of simultaneity can be understood in terms of dynamical consideration. The time-offset relation can be interpreted as the accumulated time difference between two spatially separated clocks because of pseudo-gravity experienced by the accelerated observer. This suggests that the clock at higher gravitational potential (placed along the direction of acceleration) runs faster. For the pseudo-gravitational potential difference between two points along the $x_{n}$-axis, we get $\Delta\phi=\phi_{1}-\phi_{2}=(\partial\phi/\partial x_{n})[x_{n}(1)-x_{n}(2)]$. When the system $S$ starts moving with constant velocity the gradient of potential in the system $S_{n}$ is zero. For accumulated time difference between two spatially separated clocks, we have $t^{(p)}(1)-t^{(p)}(2)=-g[x_{n}(1)-x_{n}(2)]v/(c^{2}|g|)=v[x_{n}(1)-x_{n}(2)]/c^{2}$.
@@
The next thing we must investigate is the origin of the Lorentz deformation. Here we may point out that the Lorentz deformation (changes of the relative position in the $x_{n}$ direction) is closely related to the shift in the moments of the acceleration. With the help of time-offset relation, integrating with respect to the infinitesimal value $dv$ we find $dl_{n}=dx_{n}[1+v^{2}/(2c^{2})]$, which determines the spatial geometry in the accelerated frame. This equation relates the length $dl=dx_{n}$ of the rod in the inertial frame $S$ to the length $dl_{n}$ in the accelerated frame $S_{n}$. Our analysis based on using standard measuring rods in an accelerated frame to measure it geometrical properties shows that coordinate distance $dx_{n}$ has length $dx_{n}[1+v^{2}/(2c^{2})]$. This is a purely kinematic effect, involving no forces. We conclude that the various relativistic effects are not independent of each other, but that these effects can all be understood through a general equivalence principle.
@@
We have now all the quantities we wanted. Let us put them all together into the expression for the new space-time variables. The new independent variables $(x_{n}^{\mathrm{(p)}},t_{n}^{\mathrm{(p)}})$ can be expressed in terms of the old independent variables $(x_{n},t_{n})$: $ct_{n}^{\mathrm{(p)}}=t_{n}[1-v^{2}/(2c^{2})]-(vx_{n}/c^{2})[1+v^{2}/(2c^{2})]$, $x_{n}^{\mathrm{(p)}}=x_{n}[1+v^{2}/(2c^{2})]$. Above we made a simplification in calculating time and space transformations by considering only low velocities. The physical principles that produced the space-time transformation were, however, made clear. The next question is: What is the transformation in the case of an arbitrary velocity? That is easy: the answer is $t_{n}^{\mathrm{(p)}}=t_{n}\sqrt{1-v^{2}/c^{2}}-(vx_{n}/c^{2})/\sqrt{1-v^{2}/c^{2}}$, $x_{n}^{\mathrm{(p)}}=x_{n}/\sqrt{1-v^{2}/c^{2}}$. Quantity $dt_{n}^{\mathrm{(p)}}$ characterizes physical time, which is independent on the choice of coordinate time. Let us note that $t_{n}$ in the expression for $t_{n}^{\mathrm{(p)}}$ is the coordinate time for the accelerated frame. Physical time determines the flow of time in a physical process. The $|dx_{n}^{\mathrm{(p)}}|$ is nothing, but the physical distance between two points of 3-space. The interval in physically measurable differentials is a diagonal quantity. The physically measurable differentials $dt_{n}^{\mathrm{(p)}}$ and $dx_{n}^{\mathrm{(p)}}$ are linear combinations of the coordinate time $dt_{n}$ and the coordinate differential $dx_{n}$, the coefficients of these combinations are built up of the components of the metric tensor. In new physical coordinates we obtain interval in the following form $ds^{2}=(cdt_{n}^{\mathrm{(p)}})^{2}$ $(dx_{n}^{\mathrm{(p)}})^{2}$ $=(1-v^{2}/c^{2})c^{2}dt^{2}_{n}-2vdx_{n}dt_{n}-dx_{n}^{2}$.
@@
The previous derivation of the Langevin metric in the accelerated frame from the pseudo-gravitation field includes one delicate point. Metric applies to physical laws, not to physical facts. We interpret the Langevin metric to mean that the law of electrodynamics is expressed by an equation that has the form Eq.(17) in the accelerated frame. By physical facts in this context we mean, for example, the aberration of light radiated by a single ”plane wave” emitter in the accelerated frame. The electrodynamics equation needs to be integrated with the initial condition for the radiation wavefront. After the inverse Galilean boost, the coordinates transform as $x_{n}=x-vt,t_{n}=t$, so that the wavefront of the emitted light beam is perpendicular to the vertical direction $z_{n}$ after the acceleration, Fig. 15. The Langevin metric, together with this initial condition, describes the aberration of light effect in the accelerated frame.